# On the rank index of projective curves of almost minimal degree

### v1

@@ 4. Monomial projections
    $X_{d}$ does not satisfy property $\operatorname{QR}(3)$; Moreover,
@@
We first prove that $\delta(X_{d},4)-\delta(X_{d},3)\geq 2$ for $d\geq 5$. Let $Q=\sum_{i=1}^{\frac{1}{2}(d^{2}-3d-2)}a_{i}Q_{i}$ where $a_{i}\in\mathbb{K}$ for $1\leq i\leq\frac{1}{2}(d^{2}-3d-2)$ and let $M_{Q}$ be the corresponding $\frac{1}{2}(d^{2}-3d-2)\times\frac{1}{2}(d^{2}-3d-2)$ symmetric matrix. We claim that there exist at least two independent linear forms in $\sqrt{I(4,M_{Q})}$.
We fix $Q_{1}=x_{0}x_{4}-x_{1}x_{3}$ and $Q_{2}=x_{0}x_{5}-x_{1}x_{4}$. Consider the $4\times 4$ submatrix of $M_{Q}$ obtained by selecting the rows and columns of $M_{Q}$ which correspond to $x_{0},x_{1},x_{3},x_{4}$, respectively:
$\begin{array}[]{c c}&\begin{array}[]{c c c c}x_{0}&x_{1}&x_{3}&x_{4}\\ \end{array}\\ \begin{array}[]{c c c c}x_{0}\\ x_{1}\\ x_{3}\\ x_{4}\end{array}&\begin{bmatrix}0&0&0&a_{1}\\ 0&0&-a_{1}&*\\ 0&-a_{1}&*&*\\ a_{1}&*&*&*\end{bmatrix}\end{array}.$
The determinant of this submatrix is $a_{1}^{4}$ which is contained in $I(4,M_{Q})$. Hence, $a_{1}\in\sqrt{I(4,M_{Q})}$. Now, consider the $4\times 4$ submatrix of $M_{Q}$ obtained by selecting the rows and columns of $M_{Q}$ which correspond to $x_{0},x_{1},x_{4},x_{5}$, respectively:
$\begin{array}[]{c c}&\begin{array}[]{c c c c}x_{0}&x_{1}&x_{4}&x_{5}\\ \end{array}\\ \begin{array}[]{c c c c}x_{0}\\ x_{1}\\ x_{4}\\ x_{5}\end{array}&\begin{bmatrix}0&0&a_{1}&a_{2}\\ 0&0&-a_{2}&*\\ a_{1}&-a_{2}&*&*\\ a_{2}&*&*&*\end{bmatrix}\end{array}.$
The determinant of this submatrix is $g\cdot a_{1}+a_{2}^{4}\in I(4,M_{Q})$ for $g\in\mathbb{K}[a_{1},\dots,a_{4}]_{3}$. Since $a_{1}\in\sqrt{I(4,M_{Q})}$, it follows that $a_{2}^{4}\in\sqrt{I(4,M_{Q})}$ $a_{2}\in\sqrt{I(4,M_{Q})}$. Therefore, $\delta(X_{5},4)-\delta(X_{5},3)\geq 2$
@@
    $Q_{6}=x_{3}x_{6}-x_{4}x_{5}=x_{3}x_{6}+x_{4}x_{5}-2x_{4}x_{5}\sim_{(6,2)}0$ by Lemma 4.2.
Thus, $\delta(\pi_{e_{2}}(C^{6}),4)-\delta(\pi_{e_{2}}(C^{6}),3)=2$
@@ 1. Case 2 :
5. $\operatorname{QR}(k)$ properties of $\pi_{p}(\mathcal{C}_{d})$ for rank $3$ point $p$.
In this section, we investigate whether the space of quadrics on the projected curve $\pi_{p}(\mathcal{C}_{d})$ can be spanned by rank 3 quadratic forms depending on points of rank $3$. We recall that $\pi_{p}(\mathcal{C}_{d})$ satisfies property $\operatorname{QR}(3)$ for points of rank at most two.
Suppose $p$ is a point in $\mathcal{C}_{d}^{3}\setminus\mathcal{C}_{d}^{2}$ for $d\geq 5$, and let $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})\subset\mathbb{P}^{d-1}$ be the image of $\mathcal{C}_{d}$ under the linear projection. In this case, $\mathcal{C}$ is not generated by quadrics due to the existence of a trisecant line $\mathbb{L}$ to $\mathcal{C}$, but $\mathcal{C}\cup\mathbb{L}$ is generated by quadratic polynomials. Let $X$ be the scheme defined by the quadrics in $\mathcal{C}$, i.e., $X=\mathcal{C}\cup\mathbb{L}$.
@@
By [IK99, Lemma 1.31 and Proposition 1.36], given the degree of the lower-degree generator of $f_{p}^{\perp}$, we have a unique general additive decomposition of $f_{p}$. Moreover, each term in this general additive decomposition corresponds to a point on $C_{d}$ (with multiplicities). For example, in the case of (2), from the additive decomposition of $f_{p}$, the point $p$ is spanned by a double point and a simple point on $C_{d}$ (corresponding to $\ell_{1}^{2}$ and $\ell_{2}$, respectively). Therefore, the trisecant line $\mathbb{L}$ must intersect $\pi_{p}(C_{d})$ at the image of the double point and the simple point.
@@
We first note that it suffices to consider the point $p\in\mathbb{P}^{d}$ corresponding to $f_{p}\in\mathbb{K}[s,t]_{d}$ such that $(f_{p})^{\perp}=(T^{3},g(S,T))\subset\mathbb{K}[S,T]$ by using an automorphism of $\mathbb{P}^{1}$ that sends the linear form $\ell$ to $T$ in $\mathbb{K}[S,T]_{1}$. Moreover, we can take any higher degree generator of $(f_{p})^{\perp}$ among any forms in $\{g(S,T)-T^{3}h(S,T)\mid h(S,T)\in\mathbb{K}[S,T]_{d-3}\}$. Therefore, we can assume
@@
We claim that $\sqrt{I(4,M_{Q})}$ contains at least one linear form. Following the proof of Corollary 3.2, $I(X)$ is generated by the quadratic generators of
$I(\mathcal{S}(1,d-2))=\left(2\times 2\text{ minors of }\begin{bmatrix}z_{0}&z_{2}&z_{3}&\cdots&z_{d-2}\\ z_{1}&z_{3}&z_{4}&\cdots&z_{d-1}\end{bmatrix}\right)$
@@
Let $Q=\sum_{i=0}^{N}a_{i}Q_{i}$ be the general element in $\mathbb{K}[a_{0},\dots,a_{N}]$ where $Q_{i}$ are the quadratic generators of $I(X)$ for $30\leq i\leq N=\binom{d-1}{2}-$ 3. Without loss of generality, suppose that $Q_{0}=-z_{1}z_{2}+z_{0}z_{3}$ and $Q_{1}=-z_{1}z_{3}+z_{0}z_{4}$.
@@
Hence, $X$ does not satisfy property $\operatorname{QR}(3)$ and $\delta(X,4)-\delta(X,3)\geq 2$. ∎
Considering the analysis of the case $p=e_{2}$ (see Proposition 4.6), we conjecture that
@@
For $d\leq 6$, we have observed that $\delta(X,4)-\delta(X,3)=2$ for randomly generated values of $\alpha_{0},\alpha_{1},\alpha_{2}$ provided by the computer algebra system [GS].
@@
Let $p$ be a point in $\mathbb{P}^{d}$ such that $(f_{p})^{\perp}=(\ell_{1}^{2}\ell_{2},g)\subset\mathbb{K}[S,T]$ with $\deg l_{1}=\deg l_{2}=1$ and $\deg g=d-1$. The scheme $X$ does not satisfy property $\operatorname{QR}(3)$ for $d\geq 5$.
@@
Moreover, we can take any higher-degree generator of $(f_{p})^{\perp}$ among any forms in $\{g(S,T)-T^{2}(S+\alpha T)h(S,T)\mid h(S,T)\in\mathbb{K}[S,T]_{d-3}\}$. Therefore, we can assume
@@
We claim that $\sqrt{I(4,M_{Q})}$ contains at least one linear form. Following the proof of Corollary 3.2, $I(\mathcal{C}_{d})_{2}$ is spanned by the generators of $I(S(1,d-2))$ and
@@
Then, $I(4,M_{Q})$ contains $a^{4}$ where $a$ is the coefficient of the quadratic form $z_{0}z_{3}-z_{1}z_{2}$ in $Q$ since $M_{Q}$ contains a $4\times 4$ lower triangular submatrix whose diagonal entries are $\{a,-a,-a,a\}$. Thus, $\sqrt{I(4,M_{Q})}$ contains a linear form $a$, and so $\mathcal{C}_{d}$ fails to satisfy property $\operatorname{QR}(3)$. ∎
For $d=5$, we observe that $\delta(X,4)-\delta(X,3)=1$ for randomly generated values of $\alpha,\beta_{0},\beta_{1},\beta_{2}$ provided by the computer algebra system [GS]. In addition, for a point $p=e_{1}+e_{6}\in\mathbb{P}^{6}$ (where $f_{p}^{\perp}=(S^{2}T,6S^{5}-T^{5})$), we can check that $\delta(X,4)-\delta(X,3)=1$ by [GS]. These examples support 1.4.
@@
For $d=5$, we check that $X$ satisfies property $\operatorname{QR}(3)$ for randomly generated values of $\alpha_{1},\alpha_{2},\beta_{0},\beta_{1},\beta_{2}$ using the computer algebra system [GS].
@@
Let $LT:\mathbb{P}^{6}\to\mathbb{P}^{6}$ be a linear transformation that maps a point $p$ to $e_{0}$, and let $T^{*}$ denote the $k$-linear homomorphism induced by $LT$. Let $I=(g_{1},\dots,g_{8})$ be the saturated ideal of the rational normal curve $C_{6}$, and let $J=(h_{1},\dots,h_{8})$ where each $h_{i}$ for $i=1,\dots,8$ is the quadratic form obtained by applying $T^{*}$ to the variables in $g_{i}$.
@@
where $\mathbb{K}[y_{0},y_{1},\dots,y_{d}]$ is the coordinate ring of the image under $LT$. The ideal of $X=\pi_{p}(\mathcal{C}_{d})\cup\mathbb{L}$ is given by $\left(J\cap\mathbb{K}[y_{1},\dots,y_{d}]\right)_{2}$. Considering the isomorphism $f:H^{0}(C_{6},\mathcal{O}_{\mathbb{P}^{1}}(6))\to R_{1}$, the elimination of a variable is equivalent to requiring that the coefficient of the term mapping to $y_{0}$ via the $Q_{A,B}$-map is zero given a decomposition of the line bundle $\mathcal{O}_{\mathbb{P}^{1}}(6)=A^{2}\otimes B$ on $\mathcal{C}_{6}$.
We fix the decomposition $\mathcal{O}_{\mathbb{P}^{1}}(6)=\mathcal{O}_{\mathbb{P}^{1}}(2)^{2}\otimes\mathcal{O}_{\mathbb{P}^{1}}(2)$ and let $l,m,\text{ and }n$ be general members of $\mathcal{O}_{\mathbb{P}^{1}}(2)\simeq\mathbb{K}[S,T]_{2}$. The vanishing condition of the coefficient can be decomposed into following three cases:
@@
Since we obtain degree $6$ binary forms in $\gamma_{0}$ and $\gamma_{1}$ from $Q_{\mathcal{O}_{\mathbb{P}^{1}}(2),\mathcal{O}_{\mathbb{P}^{1}}(2)}$, there are $7$ generators and there are $5$ independent quadratic forms in $I(\pi_{p}(\mathcal{C}_{6}))_{2}$ among them. Therefore, we obtain $5$ independent rank $3$ quadratic generators of $I(\pi_{p}(\mathcal{C}_{6}))$ for each cases and obtained $8$ independent generators in total, as verified by the computer software system [GS].
@@
We fix the decomposition $\mathcal{O}_{\mathbb{P}^{1}}(7)=\mathcal{O}_{\mathbb{P}^{1}}(2)^{2}\otimes\mathcal{O}_{\mathbb{P}^{1}}(3)$. Let $l$ and $m$ be general elements of $\mathcal{O}_{\mathbb{P}^{1}}(2)\simeq\mathbb{K}[S,T]_{2}$, and $n$ be a general element of $\mathcal{O}_{\mathbb{P}^{1}}(3)\simeq\mathbb{K}[S,T]_{3}$. The cases where $l$, $m$, and $n$ satisfy the vanishing condition of the coefficient are distinguished as the following three cases:
@@
We obtain $9$ independent quadratic forms in $I(\pi_{p}(\mathcal{C}_{7}))_{2}$ in each case and the 27 quadratic forms span the space $I(\pi_{p}(\mathcal{C}_{d}))_{2}$ which is verified using the computer algebra system [GS].
We now prove the claim that $X=\pi_{p}(\mathcal{C}_{d})\cup\mathbb{L}$ satisfies property $\operatorname{QR}(3)$ for the point $p=\sum_{i=0}^{d}(-1)^{i}e_{i}+(-1)^{d}\sum_{i=0}^{d}e_{i}-2e_{d}$.
@@
Let $p$ be a point in $\operatorname{\mathbb{P}}^{d}$ (for $d\geq 6$) corresponding a binary form $f_{p}$ such that $(f_{p})^{\perp}=(S^{3}-ST^{2},T^{d-1})$. Then, the scheme $X=\pi_{p}(\mathcal{C}_{d})\cup\mathbb{L}$ satisfies $\operatorname{QR}(3)$ property.
@@
where $J$ is $2\times 2$ minors of
@@
$I(X)=$ (2×2 minors of [$(2\times 2\text{ minors of }\begin{bmatrix}y_{2}&y_{3}&\cdots&y_{d-2}&y_{d-1}+\cdots+y_{1}\\ y_{3}&y_{4}&\cdots&y_{d-1}&y_{d}\end{bmatrix})$
@@
First, the cases $d=6,7$ are covered by Lemma 5.4.(1) and Lemma 5.4.(2).
@@
for $22\leq i\leq d-$ 2. Note that we didn’t check $y_{2}y_{d}-y_{3}(y_{d-1}+\cdots+y_{1})\sim_{I}0$ yet.
@@
for $52\leq i\leq d-$ 5.
Moreover, since $\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j}T^{2})=y_{3}y_{i+4}-(y_{5}+y_{3}+y_{1})y_{i+2}$ for $52\leq i\leq d-$ 5 and $y_{3}y_{i+4}-y_{5}y_{i+2}=(y_{3}y_{i+4}-y_{4}y_{i+3})+(y_{4}y_{i+3}-y_{5}y_{i+2})\sim_{I}0$, we obtain
@@
for $52\leq i\leq d-$ 5. This implies that
@@
$\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j}T)=y_{i+1}y_{d}-y_{i+2}(y_{d-1}+\cdots+y_{0})\in\mathbb{K}[z_{1},z_{2},\dots,z_{d-1}]$
for $22\leq i\leq d-$ 2.
On the other hand, by the inductive hypothesis, $w_{0}w_{i+2}-(w_{2}+w_{0})w_{i}=\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j})\in\mathbb{K}[w_{0},\dots,w_{d-2}]$ for $52\leq i\leq d-$ 5. Then,
@@
Let $k$ be the integer such that $d=2k+1$. We have
@@
=\;= $-2y_{3}(y_{d-1}+\cdots+y_{4})+2y_{5}(y_{d-1}+\cdots+y_{2})-2y_{k+2}y_{k+3}$
=\;= $-2\left((y_{3}y_{d-1}-y_{5}y_{d-3})+(y_{3}y_{d-3}-y_{5}y_{d-5})+\cdots+(y_{3}y_{4}-y_{5}y_{2})\right)$
@@
[…]

### v2

@@ 99 $x_{4}x_{5}$
    $X_{d}$ does not satisfy property $\operatorname{QR}(3)$; moreover, in the notation introduced in ˜2.6, we have
@@
We first prove that $\delta(X_{d},4)-\delta(X_{d},3)\geq 2$ for $d\geq 5$. Let $Q=\sum_{i=0}^{\frac{1}{2}(d^{2}-3d-4)}a_{i}Q_{i}$ where $a_{i}\in\mathbb{K}$ for $0\leq i\leq\frac{1}{2}(d^{2}-3d-4)$ and let $M_{Q}$ be the corresponding $d\times d$ symmetric matrix. We claim that there exist at least two independent linear forms in $\sqrt{I(4,M_{Q})}$.
We fix $Q_{0}=x_{0}x_{4}-x_{1}x_{3}$ and $Q_{1}=x_{0}x_{5}-x_{1}x_{4}$. Consider the $4\times 4$ submatrix of $M_{Q}$ obtained by selecting the rows and columns of $M_{Q}$ which correspond to $x_{0},x_{1},x_{3},x_{4}$, respectively:
$\begin{array}[]{c c}&\begin{array}[]{c c c c}x_{0}&x_{1}&x_{3}&x_{4}\\ \end{array}\\ \begin{array}[]{c c c c}x_{0}\\ x_{1}\\ x_{3}\\ x_{4}\end{array}&\begin{bmatrix}0&0&0&a_{0}\\ 0&0&-a_{0}&*\\ 0&-a_{0}&*&*\\ a_{0}&*&*&*\end{bmatrix}\end{array}$.
The determinant of this submatrix is $a_{0}^{4}$ which is contained in $I(4,M_{Q})$. Hence, $a_{0}\in\sqrt{I(4,M_{Q})}$. Now, consider the $4\times 4$ submatrix of $M_{Q}$ obtained by selecting the rows and columns of $M_{Q}$ which correspond to $x_{0},x_{1},x_{4},x_{5}$, respectively:
$\begin{array}[]{c c}&\begin{array}[]{c c c c}x_{0}&x_{1}&x_{4}&x_{5}\\ \end{array}\\ \begin{array}[]{c c c c}x_{0}\\ x_{1}\\ x_{4}\\ x_{5}\end{array}&\begin{bmatrix}0&0&a_{0}&a_{1}\\ 0&0&-a_{1}&*\\ a_{0}&-a_{1}&*&*\\ a_{1}&*&*&*\end{bmatrix}\end{array}$.
The determinant of this submatrix is $g\cdot a_{0}+a_{1}^{4}\in I(4,M_{Q})$ for a cubic polynomial $g$. Since $a_{0}\in\sqrt{I(4,M_{Q})}$, it follows that $a_{1}\in\sqrt{I(4,M_{Q})}$. Therefore, by ˜2.6, $\delta(X_{5},4)-\delta(X_{5},3)\geq 2$.
@@
    $Q_{6}=x_{3}x_{6}-x_{4}x_{5}=x_{3}x_{6}+x_{4}x_{5}-2x_{4}x_{5}\sim_{(6,2)}0$ by Lemma˜4.2.
Thus, by ˜2.6 we have $\delta(X_{6},4)-\delta(X_{6},3)=2$.
@@ 1. Case 2 :
5. $\operatorname{QR}(k)$ properties of $\pi_{p}(\mathcal{C}_{d})$ for a rank $3$ point $p$.
In this section, we investigate whether the space of quadrics on the projected curve $\pi_{p}(\mathcal{C}_{d})$ can be spanned by rank 3 quadratic forms, provided the projection center $p$ is of rank 3. We recall that $\pi_{p}(\mathcal{C}_{d})$ satisfies property $\operatorname{QR}(3)$, for points $p$ of rank at most two.
Let $\operatorname{rank}_{\mathcal{C}_{d}}(p)=l$. Then $\mathcal{C}$ is contained in S(l2,dl)S(l-2,d-l) (cf. [14, Theorem 1.1.(1)]). If $d=4$, then $l=3$ and $\mathcal{C}$ is contained in $S(1,1)$. In this case, $\mathcal{C}$ satisfies only one quadratic equation, namely that of $S(1,1)$, and hence has infinitely many trisecant lines. Therefore, we may assume $d\geq 5$.
Lemma 5.1.
Suppose $p\in\mathcal{C}_{d}^{3}\setminus\mathcal{C}_{d}^{2}$ with $d\geq 5$, and let $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})\subset\mathbb{P}^{d-1}$ be the image of $\mathcal{C}_{d}$ under the linear projection centered at $p$.
  1. (1)
    There exists the unique trisecant line $\mathbb{L}:=S(1)$ of $\mathcal{C}$.
  2. (2)
    $I(\mathcal{C}\cup\mathbb{L})$ is generated by $I(\mathcal{C})_{2}$.
Proof.
  1. (1)
    If $d\geq 5$, then $\mathcal{C}$ is contained in S(1,d3)S(1,d-3) and $\mathcal{C}\equiv H+2F$. Hence $S(1)\equiv H-(d-3)F$ is a trisecant line to $\mathcal{C}$. Conversely, if there exists a trisecant line, then it must be contained in S(1,d3)S(1,d-3), since the ideal of S(1,d3)S(1,d-3) is generated by quadrics. Consequently, $\mathbb{L}=S(1)$ is the unique trisecant line of $\mathcal{C}$.
  2. (2)
    Remark that
    $\mathcal{C}\cup\mathbb{L}\;\equiv\;2H+(5-d)F.$
    1. (a)
      By Theorem 4.3.(1) in [15], $\mathcal{C}\cup\mathbb{L}$ is aCM. Then, by Proposition 3.4 in [15], we know that $I(\mathcal{C}\cup\mathbb{L})$ is generated by quadrics.
    2. (b)
      Clearly, $I(\mathcal{C}\cup\mathbb{L})_{2}\subseteq I(\mathcal{C})_{2}$. On the other hand, $I(\mathcal{C})_{2}\subseteq I(\mathcal{C}\cup\mathbb{L})_{2}$ since $\mathbb{L}$ is the trisecant line to $\mathcal{C}$.
    In consequence, we obtain $I(\mathcal{C})_{2}=I(\mathcal{C}\cup\mathbb{L})_{2}$, and hence $I(\mathcal{C}\cup\mathbb{L})$ is generated by $I(\mathcal{C})_{2}$.
Let $X$ be the scheme defined by the quadrics in $\mathcal{C}$, i.e., $X=\mathcal{C}\cup\mathbb{L}$.
@@
By [10, Lemma 1.31 and Proposition 1.36], given the degree of the lower-degree generator of $f_{p}^{\perp}$, we have a unique general additive decomposition of $f_{p}$. Moreover, each term in this general additive decomposition corresponds to a point on $C_{d}$ (with multiplicities). For example, in the case of (2), from the additive decomposition of $f_{p}$, the point $p$ is spanned by a double point and a simple point on $C_{d}$ (corresponding to $\ell_{1}^{2}$ and $\ell_{2}$, respectively). Therefore, the trisecant line $\mathbb{L}$ must intersect $\pi_{p}(C_{d})$ at the image of the double point and the simple point.
@@
We first note that it suffices to consider the point $p\in\mathbb{P}^{d}$ corresponding to $f_{p}\in\mathbb{K}[s,t]_{d}$ such that $(f_{p})^{\perp}=(T^{3},g(S,T))\subset\mathbb{K}[S,T]$ by using an automorphism of $\mathbb{P}^{1}$ that sends the linear form $\ell$ to $T$ in $\mathbb{K}[S,T]_{1}$. Moreover, we can take any higher degree generator of $(f_{p})^{\perp}$ among any forms in $\{g(S,T)-T^{3}h(S,T)\mid h(S,T)\in\mathbb{K}[S,T]_{d-4}\}$. Therefore, we can assume
@@
We claim that $\sqrt{I(4,M_{Q})}$ contains at least one linear form. Following the proof of Corollary˜3.2, $I(X)$ is generated by the quadratic generators of
[…]
@@
Let $Q=\sum_{i=0}^{N}a_{i}Q_{i}$ be the general element in $\mathbb{K}[a_{0},\dots,a_{N}]$ where $Q_{i}$ are the quadratic generators of $I(X)$ for $0\leq i\leq N=\binom{d-1}{2}-3$. Without loss of generality, suppose that $Q_{0}=-z_{1}z_{2}+z_{0}z_{3}$ and $Q_{1}=-z_{1}z_{3}+z_{0}z_{4}$.
@@
Hence, $X$ does not satisfy property $\operatorname{QR}(3)$ and $\delta(X,4)-\delta(X,3)\geq 2$ by ˜2.6. ∎
Considering the analysis of the case $p=e_{2}$ (see Proposition˜4.6), we conjecture that
@@
For $d\leq 6$, we have observed that $\delta(X,4)-\delta(X,3)=2$ for randomly generated values of $\alpha_{0},\alpha_{1},\alpha_{2}$ provided by the computer algebra system [6].
@@
Let $p$ be a point in $\mathbb{P}^{d}$ such that $(f_{p})^{\perp}=(\ell_{1}^{2}\ell_{2},g)\subset\mathbb{K}[S,T]$ with $\deg l_{1}=\deg l_{2}=1$, $l_{1}$ and $l_{2}$ linearly independent, and $\deg g=d-1$. The scheme $X$ does not satisfy property $\operatorname{QR}(3)$ for $d\geq 5$.
@@
Moreover, we can take any higher-degree generator of $(f_{p})^{\perp}$ among all forms in $\{g(S,T)-T^{2}(S+\alpha T)h(S,T)\mid h(S,T)\in\mathbb{K}[S,T]_{d-4}\}$. Therefore, we can assume
@@
We claim that $\sqrt{I(4,M_{Q})}$ contains at least one linear form. Following the proof of Corollary˜3.2, $I(X)_{2}$ is spanned by the generators of I(S(1,d3))I(S(1,d-3)) and
@@
Then, $I(4,M_{Q})$ contains $a^{4}$ where $a$ is the coefficient of the quadratic form $z_{0}z_{3}-z_{1}z_{2}$ in $Q$ since $M_{Q}$ contains a $4\times 4$ lower triangular submatrix whose diagonal entries are $\{a,-a,-a,a\}$. Thus, $\sqrt{I(4,M_{Q})}$ contains a linear form $a$, and so $X$ fails to satisfy property $\operatorname{QR}(3)$. ∎
For $d=5$, we observe that $\delta(X,4)-\delta(X,3)=1$ for randomly generated values of $\alpha,\beta_{0},\beta_{1},\beta_{2}$ provided by the computer algebra system [6]. In addition, for a point $p=e_{1}+e_{6}\in\mathbb{P}^{6}$ (where $f_{p}^{\perp}=(S^{2}T,6S^{5}-T^{5})$), we can check that $\delta(X,4)-\delta(X,3)=1$ by [6]. These examples support ˜1.4.
@@
For $d=5$, we check that $X$ satisfies property $\operatorname{QR}(3)$ for randomly generated values of $\alpha_{1},\alpha_{2},\beta_{0},\beta_{1},\beta_{2}$ using the computer algebra system [6].
@@
Let $LT:\mathbb{P}^{6}\to\mathbb{P}^{6}$ be a linear transformation that maps a point $p$ to $e_{0}$, and let $T^{*}$ denote the $\mathbb{K}$-linear homomorphism induced by $LT$. Let $I=(g_{1},\dots,g_{8})$ be the saturated ideal of the rational normal curve $C_{6}$, and let $J=(h_{1},\dots,h_{8})$ where each $h_{i}$ for $i=1,\dots,8$ is the quadratic form obtained by applying $T^{*}$ to the variables in $g_{i}$.
@@
where $\mathbb{K}[y_{0},y_{1},\dots,y_{6}]$ is the coordinate ring of the image under $LT$. The ideal of $X=\pi_{p}(\mathcal{C}_{6})\cup\mathbb{L}$ is given by $\left(J\cap\mathbb{K}[y_{1},\dots,y_{6}]\right)_{2}$. Considering the isomorphism $f:H^{0}(C_{6},\mathcal{O}_{\mathbb{P}^{1}}(6))\to R_{1}$, the elimination of a variable is equivalent to requiring that the coefficient of the term mapping to $y_{0}$ via the $Q_{A,B}$-map is zero, given a decomposition of the line bundle $\mathcal{O}_{\mathbb{P}^{1}}(6)=A^{2}\otimes B$ on $\mathcal{C}_{6}$.
We fix the decomposition $\mathcal{O}_{\mathbb{P}^{1}}(6)=\mathcal{O}_{\mathbb{P}^{1}}(2)^{2}\otimes\mathcal{O}_{\mathbb{P}^{1}}(2)$ and let $l,m,\text{ and }n$ be general members of $H^{0}(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(2))\cong\mathbb{K}[S,T]_{2}$. The vanishing condition of the coefficient can be decomposed into following three cases:
@@
Since we obtain degree $6$ binary forms in $\gamma_{0}$ and $\gamma_{1}$ from $Q_{\mathcal{O}_{\mathbb{P}^{1}}(2),\mathcal{O}_{\mathbb{P}^{1}}(2)}$, there are $7$ generators and there are $5$ independent quadratic forms in $I(\pi_{p}(\mathcal{C}_{6}))_{2}$ among them. Therefore, we obtain $5$ independent rank $3$ quadratic generators of $I(\pi_{p}(\mathcal{C}_{6}))$ for each cases and get $8$ independent generators in total, as verified by the computer software system [6].
@@
We fix the decomposition $\mathcal{O}_{\mathbb{P}^{1}}(7)=\mathcal{O}_{\mathbb{P}^{1}}(2)^{2}\otimes\mathcal{O}_{\mathbb{P}^{1}}(3)$. Let $l$ and $m$ be general elements of $H^{0}(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(2))\cong\mathbb{K}[S,T]_{2}$, and let $n$ be a general element of $H^{0}(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(3))\cong\mathbb{K}[S,T]_{3}$. The cases where $l$, $m$, and $n$ satisfy the vanishing condition for the coefficients are distinguished to be the following three cases:
@@
We obtain $9$ independent quadratic forms in $I(\pi_{p}(\mathcal{C}_{7}))_{2}$ in each case and the 27 quadratic forms span the space $I(\pi_{p}(\mathcal{C}_{d}))_{2}$ which is verified using the computer algebra system [6].
We now prove the claim that $X=\pi_{p}(\mathcal{C}_{d})\cup\mathbb{L}$ satisfies property $\operatorname{QR}(3)$ for the point $p=\sum_{i=0}^{d}(-1)^{i}e_{i}+(-1)^{d}\sum_{i=0}^{d}e_{i}+(-1)^{d-1}2e_{d}$.
@@
Let $p$ be a point in $\operatorname{\mathbb{P}}^{d}$ (for $d\geq 6$) corresponding to a binary form $f_{p}$ such that $(f_{p})^{\perp}=(S^{3}-ST^{2},T^{d-1})$. Then, the scheme $X=\pi_{p}(\mathcal{C}_{d})\cup\mathbb{L}$ satisfies property $\operatorname{QR}(3)$.
@@
where $J$ is the ideal generated by the $2\times 2$ minors of
@@
[…]
@@
The cases $d=6,7$ are covered by Lemma˜5.5.(1) and Lemma˜5.5.(2).
@@
for $2\leq i\leq d-2$. Note that we didn’t check $y_{2}y_{d}-y_{3}(y_{d-1}+\cdots+y_{1})\sim_{I}0$ yet.
@@
for $2\leq i\leq d-5$.
Moreover, since $\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j}T^{2})=y_{3}y_{i+4}-(y_{5}+y_{3}+y_{1})y_{i+2}$ for $2\leq i\leq d-5$ and $y_{3}y_{i+4}-y_{5}y_{i+2}=(y_{3}y_{i+4}-y_{4}y_{i+3})+(y_{4}y_{i+3}-y_{5}y_{i+2})\sim_{I}0$, we obtain
@@
for $2\leq i\leq d-5$. This implies that
@@
$\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j}T)=y_{i+1}y_{d}-y_{i+2}(y_{d-1}+\cdots+y_{0})\in\mathbb{K}[y_{0},y_{1},\ldots,y_{d}]$
for $2\leq i\leq d-2$.
On the other hand, by the inductive hypothesis, $w_{0}w_{i+2}-(w_{2}+w_{0})w_{i}=\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j})\in\mathbb{K}[w_{0},\dots,w_{d-2}]$ for $2\leq i\leq d-5$. Then,
@@
Let $k$ be the integer with $d=2k+1$. We have
@@
$=\;$ $-2y_{3}(y_{d-1}+\cdots+y_{4})+2y_{5}(y_{d-1}+\cdots+y_{2})-2y_{k+2}y_{k+3}$
$=\;$ $-2\left((y_{3}y_{d-1}-y_{5}y_{d-3})+(y_{3}y_{d-3}-y_{5}y_{d-5})+\cdots+(y_{3}y_{4}-y_{5}y_{2})\right)$
@@
$\sim_{I}$  $\;0.$
Acknowledgement
We greatly appreciate the referees’ careful reading of the manuscript and their detailed and insightful suggestions. The first named author was supported by the Institute for Basic Science (IBS-R032-D1-2024-a00) and was also supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (No. RS-2024-00414849). The second named author was supported by a KIAS Individual Grant(MG083101) at Korea Institute for Advanced Study and also was supported by the Sungshin Women’s University Research Grant of 2025(H20250019). The third named author was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government (MSIT) (No. 2022R1A2C1002784).
Data availability
No data was used for the research described in the article.